# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

### v4

@@ Appendix G ANOVA Test Results
  • Reference Generation task: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001).
  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).

### v5

@@ Appendix G ANOVA Test Results
  • Reference Generation task: Claude-3.5-Sonnet (p¡.0001), DeepSeek-V3(p¡.0001), GPT-4o(p¡.0001), Qwen-2.5-72B (p¡.0001), Llama-3.2-3B (p=0.065).
  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p¡.0001), DeepSeek-V3(p¡0.05), GPT-4o(p¡.01), Qwen2.5-72B (p¡.001), Llama-3.2-3B (p¡.0001).
  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p¡.0001), DeepSeek-V3(p¡.0001), GPT-4o(p¡.0001), Qwen-2.5-72B(p¡.0001), Llama-3.2-3B(p¡.001).
  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p¡0.001).